OpenStudy (yunqin):

A line passes through (-3, -2) and is perpendicular to 3x - 2y = 7. What is the equation of the line in slope-intercept form? a) y=3/2x B) y=3/2x+4 c) y=2/3x-4 d) y=-2/3x

OpenStudy (welshfella):

first find the slope of the line 3x - 2y = 7 can you do that?

OpenStudy (yunqin):

yea

OpenStudy (welshfella):

ok so can you show me please?

OpenStudy (yunqin):

its 3 rite

OpenStudy (yunqin):

no wait ok 3x-2y=7 thats going to Be y=mx+b so the slope is m

OpenStudy (yunqin):

so its gone be y=3/2x-7/2

OpenStudy (yunqin):

its A

OpenStudy (welshfella):

3/2 is the slope of the given line but the line they require is perpendicular to the given line so what is the slope of the line they require?

OpenStudy (welshfella):

do you know this relationship ? - if m = slope of a line then slope of a perpendicular line is -1 /m

OpenStudy (yunqin):

i do now

OpenStudy (welshfella):

yea lol

OpenStudy (welshfella):

so the slope of required line = -1 / (3/2) = -2/3

OpenStudy (welshfella):

so you can now write it , in slope intercept form, as y = (-2/3) x + b but you know it passes through the point x = -3 and y = -2 so plug these into this equation and solve for b. knowing b you can write the required equation.

OpenStudy (yunqin):

so it is A

OpenStudy (welshfella):

How can it be A if the slope is -2/3???

OpenStudy (welshfella):

The equations with a slope of -2/3 are c and d.

OpenStudy (welshfella):

solve the equation for b and you'll get the answer Can you do that?

OpenStudy (welshfella):

x = -3 and y = -2 for the required equation so -2 = (-2/3)* -3 + b what is b?

OpenStudy (yunqin):

ooooooooooo B=5/3

OpenStudy (welshfella):

No (-2/3)* -3 = 6/3 = 2 so -2 = 2 + b so b = ?

OpenStudy (welshfella):

subtract 2 from both sides

OpenStudy (yunqin):

4